OpenStudy (anonymous):

convert the following exponential equations to logarithmic functions: e^x=7 and 8^x=51

OpenStudy (anonymous):

\[\ln(7)=x\]

OpenStudy (anonymous):

\\[log_8(51)=x\]

OpenStudy (anonymous):

unless you are supposed to use the change of base formula in which case the last on is \[\frac{ln(51)}{\ln(8)}=x\]

OpenStudy (anonymous):

not sure it you have to change the base

OpenStudy (anonymous):

ps thank u very much

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you there?